OpenStudy (anonymous):

The population of a city has been increasing at a rate of 1.2% per year. The population was 125,000 in the year 2000. a) find the population in 2012. b) estimate the year when the population will be 200,000

OpenStudy (anonymous):

every year the difference between population is 1.2%(i.e 1.2/100) a = 125,000 d = 1.2/100 a). \[a_n = a_0 +(n-1)d\] n = 2012 - 2000 = 12years \[a_n = 125,000 +(12-1)*(1.2/100)\] b) \[a_n = a_0 +(n-1)d\] an = 200,000 a0 = 125,000 d = 1.2/100 find n

OpenStudy (anonymous):

can you help me find n??

OpenStudy (anonymous):

@sangya21

OpenStudy (anonymous):

okkk wait total population = 200,000 sum = (n/2) [2a + (n-1)d] 200,000 = (n/2)[2* 125,000 +(n-1) 1.2/100] 200,00 = 125,000n + (1.2/100)n *(n/2) -(1.2/100)*n/2 200,000 = 0.6n^2/100 +(125,000n - 0.6n/100)